My question is... A tree that was 30meters high was broken so that the broken part was 7 times the length of the part that was left standing. Find the length of each part.
Answer by flame8855(424)

let x be the length of the standing part and y be the length of the other part
x+y = 30
y=7x
8x = 30
x= 30/8=15/4 m
x=15/4
y=7x= 105/4 m
it is the same answer
15/4 = 2.75
105/4= 26.25